# Safety Filters for Black-Box Dynamical Systems by Learning Discriminating Hyperplanes

### v1

@@ Abstract
Learning-based approaches are emerging as an effective approach for safety filters for black-box dynamical systems. Existing methods have relied on certificate functions like Control Barrier Functions (CBFs) and Hamilton-Jacobi (HJ) reachability value functions. The primary motivation for our work is the recognition that ultimately, enforcing the safety constraint as a control input constraint at each state is what matters. By focusing on this constraint, we can eliminate dependence on any specific certificate function-based design. To achieve this, we define a discriminating hyperplane that shapes the half-space constraint on control input at each state, serving as a sufficient condition for safety. This concept not only generalizes over traditional safety methods but also simplifies safety filter design by eliminating dependence on specific certificate functions. We present two strategies to learn the discriminating hyperplane: (a) a supervised learning approach, using pre-verified control invariant sets for labeling, and (b) a reinforcement learning (RL) approach, which does not require such labels. The main advantage of our method, unlike conventional safe RL approaches, is the separation of performance and safety. This offers a reusable safety filter for learning new tasks, avoiding the need to retrain from scratch. As such, we believe that the new notion of the discriminating hyperplane offers a more generalizable direction towards designing safety filters, encompassing and extending existing certificate-function-based or safe RL methodologies.
@@ 1 Introduction
While learning-based control has demonstrated capabilities in solving complex tasks for uncertain dynamical systems, safety assurance remains an unresolved challenge. Safe Reinforcement Learning (RL) has made strides towards co-learning performance and safety (Garcıa and Fernández(2015)), but the absence of provable guarantees in these data-driven methods remains a significant challenge. Conversely, although model-based control theory has well-established mechanisms to enforce safety constraints, its application to uncertain systems is limited when the models deviate from the actual system dynamics.
In an effort to bridge this gap and provide safety guarantees for black-box dynamical systems with unknown closed-form expressions, researchers are actively combining model-based control techniques with data-driven methods (Brunke et al.(2022a)Brunke, Greeff, Hall, Yuan, Zhou, Panerati, and Schoellig; Wabersich et al.(2023)Wabersich, Taylor, Choi, Sreenath, Tomlin, Ames, and Zeilinger). Traditional model-based methods for safety problems fundamentally involve two key components: a safe set where the trajectory can remain indefinitely, and a control input constraint that guarantees this invariance within the set. These methods typically rely on certificate functions such as Lyapunov functions, Control Barrier Functions (Ames et al.(2016)Ames, Xu, Grizzle, and Tabuada), or HJ reachability value functions (Fisac et al.(2019a)Fisac, Akametalu, Zeilinger, Kaynama, Gillula, and Tomlin). Consequently, existing learning-based methods for designing safety filters for black-box systems have primarily focused on learning the certificate functions (Dawson et al.(2023)Dawson, Gao, and Fan).
@@
Towards this end, we define a discriminating hyperplane that shapes the half-space constraint on control input at each state and serves as a sufficient condition to ensure control invariance of a safe set. The discriminating hyperplane-based safety constraint generalizes the well-known Nagumo-like condition (Blanchini(1999)), CBF constraint (Ames et al.(2016)Ames, Xu, Grizzle, and Tabuada), and HJ reachability optimal control law (Fisac et al.(2019a)Fisac, Akametalu, Zeilinger, Kaynama, Gillula, and Tomlin), for control-affine systems.
Next, we present two approaches for learning the discriminating hyperplane for black-box systems. The first method utilizes supervised learning, where the learning is guided by labels derived from a pre-verified control invariant set. Although this assumes access to prior knowledge of the invariant set, it can be well supplemented by leveraging existing research focused on constructing or learning control invariant sets (Bertsekas(1972); Blanchini(1999); Fisac et al.(2019a)Fisac, Akametalu, Zeilinger, Kaynama, Gillula, and Tomlin; Wabersich and Zeilinger(2018); Bansal and Tomlin(2021); Cao et al.(2022)Cao, Bloch, and Coogan). In our simulations across various systems, we find that the safety filter, based on the learned discriminating hyperplane, closely approximates a filter that could be designed using ground-truth system dynamics.
The second method, utilizing reinforcement learning (RL), circumvents the need for access to control invariant sets by learning the discriminating hyperplane directly from trajectory data. We can employ existing RL algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al.(2017)Schulman, Wolski, Dhariwal, Radford, and Klimov), for this purpose. A key strength of this approach is the ability to separate a learning-based policy into two components–task achievement and safety filtering. This separation allows the learned safety filter to be reused or adapted to various performance tasks. The concept of separating performance and safety is well-supported by existing safe RL literature (Thananjeyan et al.(2021)Thananjeyan, Balakrishna, Nair, Luo, Srinivasan, Hwang, Gonzalez, Ibarz, Finn, and Goldberg; Wagener et al.(2021)Wagener, Boots, and Cheng; Kim et al.(2023b)Kim, Oh, Lee, Choi, Ji, Jung, Youm, and Hwangbo). The main distinction of our work is that there is no on-off switching mechanism between performance and safety policies. Instead, we integrate these two aspects into a unified safety filter design based on the discriminating hyperplane concept.
@@ 2 Related Work
Certificate functions: Many model-based control approaches for safety problems use the concept of certificate functions (Prajna(2006); Dawson et al.(2023)Dawson, Gao, and Fan). These are also referred to as safety index in Liu and Tomizuka(2014) or energy function in Wei and Liu(2019). Put simply, a certificate function is a state-dependent scalar function, whose level sets characterize the safe domain and whose gradient can impose a constraint on the control input to ensure safety. To address difficulties in designing certificate functions using classical methods, learning-based methods are being actively explored (Fisac et al.(2019b)Fisac, Lugovoy, Rubies-Royo, Ghosh, and Tomlin; Srinivasan et al.(2020)Srinivasan, Eysenbach, Ha, Tan, and Finn; Huh and Yang(2020); Lindemann et al.(2021)Lindemann, Robey, Jiang, Tu, and Matni; Thananjeyan et al.(2021)Thananjeyan, Balakrishna, Nair, Luo, Srinivasan, Hwang, Gonzalez, Ibarz, Finn, and Goldberg; Liu et al.(2023)Liu, Liu, and Dolan; So et al.(2023)So, Serlin, Mann, Gonzales, Rutledge, Roy, and Fan; Castañeda et al.(2023)Castañeda, Nishimura, McAllister, Sreenath, and Gaidon).
@@
Learning constraints for uncertain systems: A crucial step in designing the safety filter is to come up with a valid constraint to be imposed on the control input. Usually, this constraint is a sufficient condition for the closed-loop trajectory to stay invariant inside the safe domain. These constraints can be derived from certificate functions. For uncertain systems, these are typically designed based on a nominal model, and are combined with data-driven methods to accommodate the modeling error (Choi et al.(2020)Choi, Castañeda, Tomlin, and Sreenath; Castañeda et al.(2021)Castañeda, Choi, Zhang, Tomlin, and Sreenath; Taylor et al.(2020)Taylor, Singletary, Yue, and Ames; Taylor et al.(2021)Taylor, Dorobantu, Dean, Recht, Yue, and Ames; Brunke et al.(2022a)Brunke, Greeff, Hall, Yuan, Zhou, Panerati, and Schoellig; Wabersich et al.(2023)Wabersich, Taylor, Choi, Sreenath, Tomlin, Ames, and Zeilinger). Our methodology aligns closely with these techniques. However, crucially, our approach is independent of both certificate functions and nominal models.
@@
Safe Reinforcement Learning (RL): Numerous safe RL algorithms proposed in Garcıa and Fernández(2015); Achiam et al.(2017)Achiam, Held, Tamar, and Abbeel; Ray et al.(2019)Ray, Achiam, and Amodei; Srinivasan et al.(2020)Srinivasan, Eysenbach, Ha, Tan, and Finn; Thananjeyan et al.(2021)Thananjeyan, Balakrishna, Nair, Luo, Srinivasan, Hwang, Gonzalez, Ibarz, Finn, and Goldberg; Wagener et al.(2021)Wagener, Boots, and Cheng address safety using a purely data-driven approach. A majority of these methods construe safety violations as an accumulation of safety-relevant cost, referred to as the constrained Markov Decision Process problem (Altman(1998)). While such a formulation finds applicability in certain safety contexts, it falls short in capturing instantaneous constraint violations like collision. While some safe RL algorithms incorporate control theory concepts—for example, Lyapunov theory in Chow et al.(2018)Chow, Nachum, Duenez-Guzman, and Ghavamzadeh and reachability theory in Fisac et al.(2019b)Fisac, Lugovoy, Rubies-Royo, Ghosh, and Tomlin; Huh and Yang(2020); Hsu et al.(2023)Hsu, Nguyen, and Fisac—these ideas are applied in the algorithmic design rather than in the structural design of the learned safe policy. In contrast, our approach imposes a structural knowledge derived from the discriminating hyperplane to the learned safety filter. Approaches in Cheng et al.(2019)Cheng, Orosz, Murray, and Burdick; Emam et al.(2022)Emam, Notomista, Glotfelter, Kira, and Egerstedt explicitly use certificate functions to achieve this.
@@ 3.1 Safety filters for black-box dynamical systems
[…]
@@
The safety problems we focus on is to ensure that the system states satisfy specific constraints over an indefinite time horizon. Thus, safety for the system (1) is encoded by a target constraint set $X\subset\mathbb{R}^{n}$ that must be respected during the evolution of the system:
@@
The safety filter we aim to design operates as an intermediary between the reference controller $\pi_{\text{ref}}:\mathbb{R}^{n}\rightarrow U$ and the system, ensuring that the downstream control adheres to the safety constraint described in (2). Reference controllers are typically safety-agnostic and can take various forms, from a neural network policy optimized for performance objectives, hand-designed controllers from domain experts, to a human operator’s commands. The safety filter $\pi_{\text{safe}}:\mathbb{R}^{n}\rightarrow U$ modifies the input signal from the reference controller when necessary, yielding a control input signal $\mathrm{u}(t)=\pi_{\text{safe}}(\mathrm{x}(t);\pi_{\text{ref}})$. We mainly focus on the safety constraint within the filter, which regulates the control input to guarantee safety:
@@
We say that $c(x,u)\geq 0$ is a (valid) safety constraint if there exists a Lipschitz feedback policy $\pi:\mathbb{R}^{n}\rightarrow U$ and a safe domain $S\subseteq X$ such that the constraint $c(x,\pi(x))\geq 0$ is feasible for all $x\in S$, and the trajectory resulting from the control signal $\mathrm{u}(t)=\pi(\mathrm{x}(t))$ satisfies (2) for all $x\in S$ where $\mathrm{x}(0)=x$.
Note that the Lipschitz continuity is required to guarantee the solution existence and uniqueness of $\mathrm{x}(\cdot)$. In words, the safety constraint is a sufficient condition that, when $u$ satisfies the constraint, safety is guaranteed for all states in the safe domain $S$, a subset of the target constraint set $X$. Consequently, the effective design of a safety filter hinges on the choice of a valid safe domain $S$ and a safety constraint $c(x,u)\geq 0$.
We first review how a safe domain is generally designed in the literature. Many existing design approaches (Wabersich et al.(2023)Wabersich, Taylor, Choi, Sreenath, Tomlin, Ames, and Zeilinger) seek to find a set $S$ that is control invariant, defined next.
Definition 2 (Control Invariance (Blanchini(1999)))
@@
The definition implies that a control invariant set $S$ that is a subset of $X$ can serve as a safe domain. The representation of the control invariant set can vary significantly depending on the chosen design methodology. It ranges from geometric structures like polytopes (Blanchini(1999)) to level sets of scalar functions such as certificate functions (Dawson et al.(2023)Dawson, Gao, and Fan), and extends to the concept of feasibility in receding-horizon optimal control problems (Wabersich and Zeilinger(2021)). This versatility underscores control invariance as a foundational concept in designing safe domains.
@@ 3.2 Discriminating hyperplane for safety constraint
Once a control invariant safe domain $S$ is specified, a similar blueprint for characterizing a valid safety constraint exists. It is based on the geometric relationship that the control invariant set and the vector field of the dynamics satisfy at its boundary, $\partial S$, which is known as the “Nagumo-like” condition in the literature (Nagumo(1942); Blanchini(1999); Aubin et al.(2011)Aubin, Bayen, and Saint-Pierre; Ames et al.(2016)Ames, Xu, Grizzle, and Tabuada):
@@
(Tangential characterization of control invariant sets Aubin et al.(2011)Aubin, Bayen, and Saint-Pierre) Let the dynamics (1) satisfy Assumption 1. Then, a closed set $S\subset\mathbb{R}^{n}$ is control invariant if and only if for all $x\in\partial S$,
@@
where $h\!:\!\mathbb{R}^{n}\rightarrow\mathbb{R}$ is continuously differentiable function such that $S\!=\!\{x|h(x)\geq 0\}$ and $\frac{\partial h}{\partial x}(x)$ is bounded away from 0 for all $x\in\partial S$, implying that the slope of $h$ does not vanish at the boundary [^1] Such $h$ exists for $S$ whose interior is not empty and boundary is continuously differentiable (Lieberman(1985))..
The condition (4) is in fact a geometric condition implying that for the set to be control invariant, there must exist a control input that renders the vector field pointing inward to the set. Thus, the condition remains invariant under different choices of the distance-like function $h$, and (4) is merely a mathematical description of this geometric property. Extending this property, in the new notion of the discriminating hyperplane we introduce next, the resulting safety constraint also does not rely on any choice of $h$. In contrast, the notion of CBF in (Ames et al.(2016)Ames, Xu, Grizzle, and Tabuada), which is also inspired by Lemma 1, results in a certifying constraint that is dependent on $h$.
From (4), a constraint which is affine in $u$, imposed whenever the state $x$ is at the boundary of $S$, is sufficient to regulate the trajectory to not exit the set $S$. This serves as a starting point for constructing a control-affine constraint extended to all $x\in S$ including the interior of the set as well. By noticing that the control-affine inequality defines a halfspace in the control input space, we define the discriminating hyperplane as below:
@@
A discriminating hyperplane for systems satisfying Assumption 1 and a control invariant set $S\in\mathbb{R}^{n}$, is a hyperplane defined in the control input space represented by $a(x)^{\top}u=b(x)$, for each state $x\in S$. The resulting half-space constraint $a(x)^{\top}u\geq b(x)$ is then a safety constraint as defined in Definition 1.
In words, the discriminating hyperplane discriminates the control input between certified input that guarantees safety, and uncertified input that can potentially lead to safety violations. The hyperplane is defined in the control input space and is parameterized by each state. If such a discriminating hyperplane can be determined, we can use it effectively to construct the following safety filter:   Discriminating hyperplane-based min-norm safety filter:
$\pi_{\text{safe}}(x)=\arg\min_{u\in U}\quad$ $||u-\pi_{\text{ref}}(x)||^{2}$ (5)
@@
Next, we present a sufficient condition for $\{x\in\mathbb{R}^{n}|a(x)^{\top}u=b(x)\}$ to be a discriminating hyperplane. In words, the theorem says that any discriminating hyperplane leading to the satisfaction of the Nagumo-like condition (4) is valid.
@@
For the control input set $U$ that is polytopic, if $a:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and $b:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are Lipschitz continuous in $x$, and if for all $x\in\partial S$, $\{u\in U\;|\;a(x)^{\top}u\geq b(x)\}$ has a nonempty interior and is a subset of $\{u\in U\;|\;\frac{\partial h}{\partial x}(x)\cdot(f(x)+g(x)u)\geq 0\}$, then $a$ and $b$ define a discriminating hyperplane, $\{x\in\mathbb{R}^{n}|a(x)^{\top}u=b(x)\}$. (Proof: Appendix 7.1)
@@
[…]
@@
CBFs in (Ames et al.(2016)Ames, Xu, Grizzle, and Tabuada) and HJ reachability value functions in (Fisac et al.(2019a)Fisac, Akametalu, Zeilinger, Kaynama, Gillula, and Tomlin) offer special cases of the discriminating hyperplane. For a CBF $h$, the hyperplane is given by $a(x)=\frac{\partial h}{\partial x}(x)\cdot g(x)$ and $b(x)=-\frac{\partial h}{\partial x}(x)\cdot f(x)-\alpha(h(x))$, where $\alpha$ is the comparison function associated with the CBF. For a reachability value function $V$, the hyperplane is given by $a(x)=\frac{\partial V}{\partial x}(x)\cdot g(x)$ and $b(x)=-\frac{\partial V}{\partial x}(x)\cdot f(x)$. As such, the discriminating hyperplane can be considered as a generalized structure of safety constraints for control-affine systems that unifies and extends the results of the existing methods.
4 Learning Discriminating Hyperplane
Having defined the discriminating hyperplane and its role in shaping the safety constraint, this section presents supervised and reinforcement learning approaches to learn the discriminating hyperplane from black-box system trajectory data.
4.1 Supervised learning approach
In order to formulate a supervised learning problem, we first present a simple way to construct a discriminating hyperplane for a given control invariant set $S$, without having to rely on any specific distance-like function $h$ of the set $S$, as in (4) or for the case of CBF. This is consistent with the core principle of our approach, which is to avoid reliance on any specific form of certificate functions or safe domain representation. Thus, we only rely on a minimal knowledge of the invariant set: an indicator of whether a given state is within $S$: $I_{S}(x)=1$ if $x\in S$, and 00 otherwise.

### v2

@@ Abstract
Learning-based approaches are emerging as an effective approach for safety filters for black-box dynamical systems. Existing methods have relied on certificate functions like Control Barrier Functions (CBFs) and Hamilton-Jacobi (HJ) reachability value functions. The primary motivation for our work is the recognition that, ultimately, enforcing the safety constraint as a control input constraint at each state is what matters. By focusing on this constraint, we can eliminate dependence on any specific certificate function-based design. To achieve this, we define a discriminating hyperplane that shapes the half-space constraint on control input at each state, serving as a sufficient condition for safety. This concept not only generalizes over traditional safety methods but also simplifies safety filter design by eliminating dependence on specific certificate functions. We present two strategies to learn the discriminating hyperplane: (a) a supervised learning approach, using pre-verified control invariant sets for labeling, and (b) a reinforcement learning (RL) approach, which does not require such labels. The main advantage of our method, unlike conventional safe RL approaches, is the separation of performance and safety. This offers a reusable safety filter for learning new tasks, avoiding the need to retrain from scratch. As such, we believe that the new notion of the discriminating hyperplane offers a more generalizable direction towards designing safety filters, encompassing and extending existing certificate-function-based or safe RL methodologies.
@@ 1 Introduction
While learning-based control has demonstrated capabilities in solving complex tasks for uncertain dynamical systems, safety assurance remains an unresolved challenge. Safe Reinforcement Learning (RL) has made strides towards co-learning performance and safety (Garcıa and Fernández, 2015), but the absence of provable guarantees in these data-driven methods remains a significant challenge. Conversely, although model-based control theory has well-established mechanisms to enforce safety constraints, its application to uncertain systems is limited when the models deviate from the actual system dynamics.
In an effort to bridge this gap and provide safety guarantees for black-box dynamical systems with unknown closed-form expressions, researchers are actively combining model-based control techniques with data-driven methods (Brunke et al., 2022a; Wabersich et al., 2023). Traditional model-based methods for safety problems fundamentally involve two key components: a safe set where the trajectory can remain indefinitely, and a control input constraint that guarantees this invariance within the set. These methods typically rely on certificate functions such as Lyapunov functions, Control Barrier Functions (Ames et al., 2016), or HJ reachability value functions (Fisac et al., 2019a). Consequently, existing learning-based methods for designing safety filters for black-box systems have primarily focused on learning the certificate functions (Dawson et al., 2023).
@@
Towards this end, we define a discriminating hyperplane that shapes the half-space constraint on control input at each state and serves as a sufficient condition to ensure control invariance of a safe set. The discriminating hyperplane-based safety constraint generalizes the well-known Nagumo-like condition (Blanchini, 1999), CBF constraint (Ames et al., 2016), and HJ reachability optimal control law (Fisac et al., 2019a), for control-affine systems.
Next, we present two approaches for learning the discriminating hyperplane for black-box systems. The first method utilizes supervised learning, where the learning is guided by labels derived from a pre-verified control invariant set. Although this assumes access to prior knowledge of the invariant set, it can be well supplemented by leveraging existing research focused on constructing or learning control invariant sets (Bertsekas, 1972; Blanchini, 1999; Fisac et al., 2019a; Wabersich and Zeilinger, 2018; Bansal and Tomlin, 2021; Cao et al., 2022). In our simulations across various systems, we find that the safety filter, based on the learned discriminating hyperplane, closely approximates a filter that could be designed using ground-truth system dynamics.
The second method, utilizing reinforcement learning (RL), circumvents the need for access to control invariant sets by learning the discriminating hyperplane directly from trajectory data. We can employ existing RL algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), for this purpose. A key strength of this approach is the ability to separate a learning-based policy into two components–task achievement and safety filtering. This separation allows the learned safety filter to be reused or adapted to various performance tasks. The concept of separating performance and safety is well-supported by existing safe RL literature (Thananjeyan et al., 2021; Wagener et al., 2021; Kim et al., 2023b). The main distinction of our work is that there is no on-off switching mechanism between performance and safety policies. Instead, we integrate these two aspects into a unified safety filter design based on the discriminating hyperplane concept.
@@ 2 Related Work
Certificate functions: Many model-based control approaches for safety problems use the concept of certificate functions (Prajna, 2006; Dawson et al., 2023). These are also referred to as safety index in Liu and Tomizuka (2014) or energy function in Wei and Liu (2019). Put simply, a certificate function is a state-dependent scalar function, whose level sets characterize the safe domain and whose gradient can impose a constraint on the control input to ensure safety. To address difficulties in designing certificate functions using classical methods, learning-based methods are being actively explored (Fisac et al., 2019b; Srinivasan et al., 2020; Huh and Yang, 2020; Lindemann et al., 2021; Thananjeyan et al., 2021; Liu et al., 2023; So et al., 2023; Castañeda et al., 2023).
@@
Learning constraints for uncertain systems: A crucial step in designing the safety filter is to come up with a valid constraint to be imposed on the control input. Usually, this constraint is a sufficient condition for the closed-loop trajectory to stay invariant inside the safe domain. These constraints can be derived from certificate functions. For uncertain systems, these are typically designed based on a nominal model, and are combined with data-driven methods to accommodate the modeling error (Choi et al., 2020; Castañeda et al., 2021; Taylor et al., 2020, 2021; Brunke et al., 2022a; Wabersich et al., 2023). Our methodology aligns closely with these techniques. However, crucially, our approach is independent of both certificate functions and nominal models.
@@
Safe Reinforcement Learning (RL): Numerous safe RL algorithms proposed in Garcıa and Fernández (2015); Achiam et al. (2017); Ray et al. (2019); Srinivasan et al. (2020); Thananjeyan et al. (2021); Wagener et al. (2021) address safety using a purely data-driven approach. A majority of these methods construe safety violations as an accumulation of safety-relevant cost, referred to as the constrained Markov Decision Process problem (Altman, 1998). While such a formulation finds applicability in certain safety contexts, it falls short in capturing instantaneous constraint violations like collision. While some safe RL algorithms incorporate control theory concepts—for example, Lyapunov theory in Chow et al. (2018) and reachability theory in Fisac et al. (2019b); Huh and Yang (2020); Hsu et al. (2023)—these ideas are applied in the algorithmic design rather than in the structural design of the learned safe policy. In contrast, our approach imposes a structural knowledge derived from the discriminating hyperplane to the learned safety filter. Approaches in Cheng et al. (2019); Emam et al. (2022) explicitly use certificate functions to achieve this.
@@ 3.1 Safety filters for black-box dynamical systems
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t))+g(\mathrm{x}(t))\mathrm{u}(t)\;\;\text{for}\;\;t>0,\qquad\mathrm{x}(0)=x,\vspace{-0.25em}$ (1)
@@
The safety problem we focus on is to ensure that the system states satisfy specific constraints over an indefinite time horizon. Thus, safety for the system (1) is encoded by a target constraint set $X\subset\mathbb{R}^{n}$ that must be respected during the evolution of the system:
@@
The safety filter we aim to design operates as an intermediary between the reference controller $\pi_{\text{ref}}:\mathbb{R}^{n}\rightarrow U$ and the system, ensuring that the downstream control adheres to the safety constraint described in (2). Reference controllers are typically safety-agnostic and can take various forms, from a neural network policy optimized for performance objectives, hand-designed controllers from domain experts, to a human operator’s commands. The safety filter $\pi_{\text{safe}}:\mathbb{R}^{n}\rightarrow U$ modifies the input signal from the reference controller when necessary, yielding a control input signal $\mathrm{u}(t)=\pi_{\text{safe}}(\mathrm{x}(t);\pi_{\text{ref}})$. We mainly focus on the safety constraint within the filter, which modifies the control input to guarantee safety:
@@
We say that $c(x,u)\geq 0$ is a (valid) safety constraint if there exists a safe domain $S\subseteq X$, such that for all Lipschitz feedback policy $\pi:\mathbb{R}^{n}\rightarrow U$ that satisfies $c(x,\pi(x))\geq 0$ for all $x\in S$, the trajectory resulting from the control signal $\mathrm{u}(t)=\pi(\mathrm{x}(t))$ satisfies (2) for all $x\in S$ where $\mathrm{x}(0)=x$.
Note that the Lipschitz continuity of $\pi$ is required to guarantee the solution $\mathrm{x}(\cdot)$ exists and is unique. In words, the safety constraint is a sufficient condition that, when $u$ satisfies the constraint, safety is guaranteed for all states in the safe domain $S$, a subset of the target constraint set $X$. Consequently, the effective design of a safety filter hinges on the choice of a valid safe domain $S$ and a safety constraint $c(x,u)\geq 0$.
We first review how a safe domain is generally designed in the literature. Many existing design approaches (Wabersich et al., 2023) seek to find a set $S$ that is control invariant, defined next.
Definition 2 (Control Invariance (Blanchini, 1999))
@@
The definition implies that a control invariant set $S$ that is a subset of $X$ can serve as a safe domain. The representation of the control invariant set can vary significantly depending on the chosen design methodology. It ranges from geometric structures like polytopes (Blanchini, 1999) to level sets of scalar functions such as certificate functions (Dawson et al., 2023), and extends to the concept of feasibility in receding-horizon optimal control (Wabersich and Zeilinger, 2021). This versatility underscores control invariance as a foundational concept in designing safe domains.
@@ 3.2 Discriminating hyperplane for safety constraint
Once a control invariant safe domain $S$ is specified, a similar blueprint for characterizing a valid safety constraint exists. It is based on the geometric relationship that the control invariant set and the vector field of the dynamics satisfy at its boundary, $\partial S$, which is known as the “Nagumo-like” condition in the literature (Nagumo, 1942; Blanchini, 1999; Aubin et al., 2011; Ames et al., 2016):
@@
(Tangential characterization of control invariant sets Aubin et al. (2011, Theorem 11.3.4)) Let the dynamics (1) satisfy Assumption 1. Then, a closed set $S\subset\mathbb{R}^{n}$ is control invariant if and only if for all $x\in\partial S$,
@@
where $h\!:\!\mathbb{R}^{n}\rightarrow\mathbb{R}$ is continuously differentiable function such that $S\!=\!\{x|h(x)\geq 0\}$ and $\frac{\partial h}{\partial x}(x)$ is bounded away from 0 for all $x\in\partial S$, implying that the slope of $h$ does not vanish at the boundary [^1] Such $h$ exists for $S$ whose interior is not empty and boundary is continuously differentiable (Lieberman, 1985)..
The condition (4) implies that for the set to be control invariant, there must exist a control input that renders the vector field pointing inward to the set. In fact, the condition is a geometric property that remains invariant under different choices of the distance-like function $h$, and (4) is merely its analytic description222(4) can be rewritten as $\exists u\!\in\!U$ s.t. $(f(x)\!+\!g(x)u)\!\in\!T_{S}(x)$, where $T_{S}$ is (Bouligand’s) tangent cone to $S$ (Clarke et al., 2008).. Extending this property, in the new notion of the discriminating hyperplane we introduce next, the resulting safety constraint also does not rely on any choice of $h$. In contrast, the notion of CBF in (Ames et al., 2016), which is also inspired by Lemma 1, results in a certifying constraint that is dependent on $h$.
From (4), a constraint which is affine in $u$, imposed whenever the state $x$ is at the boundary of $S$, is sufficient to regulate the trajectory to not exit the set $S$. This serves as a starting point for constructing a control-affine constraint extended to all $x\in S$ including the interior of the set. By noticing that the control-affine inequality defines a halfspace in the control input space, we define the discriminating hyperplane as below:
@@
A discriminating hyperplane for systems satisfying Assumption 1 and a control invariant set $S\in\mathbb{R}^{n}$ is a hyperplane defined in the control input space represented by $a(x)^{\top}u=b(x)$, for each state $x\in S$, such that the resulting half-space constraint $a(x)^{\top}u\geq b(x)$ is a safety constraint according to Definition 1.
In words, the discriminating hyperplane distinguishes between certified input signals that guarantee safety, and uncertified input signals that could potentially lead to safety violations. The hyperplane is defined in the control input space and is parameterized by each state. If such a discriminating hyperplane can be determined, we can use it effectively to construct the following safety filter:   Discriminating hyperplane-based min-norm safety filter:
$\pi_{\text{safe}}(x)=\arg\min_{u\in U}$ $||u-\pi_{\text{ref}}(x)||^{2}$ (5)
@@
Next, we present a sufficient condition for $a(x)^{\top}u=b(x)$ to be a discriminating hyperplane. In words, the theorem says that any discriminating hyperplane leading to the satisfaction of the Nagumo-like condition (4) is valid.
@@
For the control input set $U$ that is polytopic, if $a:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and $b:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are Lipschitz continuous in $x$, and if for all $x\in\partial S$, $\{u\in U\;|\;a(x)^{\top}u\geq b(x)\}$ has a nonempty interior and is a subset of $\{u\in U\;|\;\frac{\partial h}{\partial x}(x)\cdot(f(x)+g(x)u)\geq 0\}$, then $a$ and $b$ define a discriminating hyperplane, $\{u\;|\;a(x)^{\top}u=b(x)\}$. (Proof: Appendix 7.1)
@@
If $\pi_{\text{ref}}:\mathbb{R}^{n}\rightarrow U$ is Lipschitz continuous in $x$ and if $a,b$ satisfy the conditions in Theorem 1, then $\pi_{\text{safe}}$ in (5) is also Lipschitz continuous in $x$.
@@
CBFs in (Ames et al., 2016) and HJ reachability value functions in (Fisac et al., 2019a) offer special cases of the discriminating hyperplane. For a CBF $h$, the hyperplane is given by $a(x)=\frac{\partial h}{\partial x}(x)\cdot g(x)$ and $b(x)=-\frac{\partial h}{\partial x}(x)\cdot f(x)-\alpha(h(x))$, where $\alpha$ is the comparison function associated with the CBF. For a reachability value function $V$, the hyperplane is given by $a(x)=\frac{\partial V}{\partial x}(x)\cdot g(x)$ and $b(x)=-\frac{\partial V}{\partial x}(x)\cdot f(x)$. As such, the discriminating hyperplane can be considered as a generalized structure of safety constraints for control-affine systems that unifies and extends the results of the existing methods.
3.3 Sample-and-hold lookahead-based discriminating hyperplane
We present a simple way to construct a discriminating hyperplane for a given control invariant set $S$, without having to rely on any specific distance-like function $h$ of the set $S$. This is consistent with the core principle of our approach, which is to avoid reliance on any specific form of certificate functions or safe domain representation. Thus, we only rely on a minimal knowledge of the invariant set: an indicator of whether a given state is within $S$: $I_{S}(x)=1$ if $x\in S$, and 00 otherwise.